OpenStudy (anonymous):

There are 48 students in a school play. The ratio of boys to girls is 5:7. How many boys are in the school play? Explain how you found your answer.

OpenStudy (anonymous):

meow

OpenStudy (anonymous):

MEOW!!!!!!!!!!!!!!!

OpenStudy (anonymous):

purrrrrrr

OpenStudy (zpupster):

boys/girls=5/7 boys=5x where x is the no of students girls=7x 5x+7x=48 12x=48 x=4 5(4)=20 boys in play 48 - boys =girls in play